Calcium hydroxide and phosphoric acid react to form calcium phosphate and water. Find the moles of water, if you have 6.50 moles of calcium hydroxide.


1
Expert's answer
2021-03-22T06:27:00-0400

Solution:

The balanced chemical equation:

3Ca(OH)2 + 2H3PO4 → Ca3(PO4)2 + 6H2O

According to the equation above:

Moles of Ca(OH)2 / 3 = Moles of H2O / 6

Moles of Ca(OH)2 = Moles of H2O / 2


Moles of H2O = 2 × Moles of Ca(OH)2 = 2 × 6.5 mol = 13 mol

Moles of H2O = 13 mol H2O


Answer: 13 mol of water (H2O).